OpenStudy (cometailcane):

The mens NCAA basketball tournament starts today with 64 teams. After each round is played, half of the teams are left. This continues until one team is left and declared champion. This format can be represented by an exponential function, where 1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you are dividing by two each time you can write it as \[60\times \left(\frac{1}{2}\right)^n\]

OpenStudy (misty1212):

or \[\frac{64}{2^n}\] or even \[64\times 2^{-n}\]

OpenStudy (cometailcane):

oh so thats the answer?

OpenStudy (misty1212):

yeah pick one i would go with the first one

OpenStudy (cometailcane):

thank you! seems so much harder than it actually is haha